# Low Mass Naked Singularities from Dark Core Collapse

### v1

@@ 7 Constraining $\sigma-m_{\chi}$ parameter space from the Missing pulsar problem
Point (1) should give a lower limit of $\sigma$ and Point (2) should give an upper limit of $\sigma$, each as a function of $m_{\chi},~{}\mathcal{D}$, DM profile, particle type (boson or fermion), etc. Thus, the $t_{\rm d}=t_{\rm Ga}$ condition should ideally give a value of $\sigma$ as a function of $m_{\chi},~{}\mathcal{D}$, DM profile, particle type (boson or fermion), etc. The transmutation timescale $t_{\rm d}$ can be calculated using Eq. (30) and Eq. (14).
(a)  NFW DM density profile (Eq. 31): $\rho_{\rm DM}=800$ GeV.cm-3
(b)  Cored DM profile (Eq. 34): $\rho_{\rm DM}=17$ GeV.cm-3
(c)  Cored DM profile (Eq. 34): $\rho_{\rm DM}=410$ GeV.cm-3
@@
Figure 4: Constraints on the DM-nucleon scattering cross section $\sigma$, as a function of the PDM particle mass $m_{\chi}$ for the possible transmutation. The red-solid / blue-dashed curve corresponds to $t_{\rm d}=t_{\rm Ga}$ for boson/fermion, respectively, assuming all the old pulsars located within a distance $\mathcal{D}\sim 10$ pc around the Galactic center have been transmuted into BH or naked singularity. The parameter space located on the right side of the red-solid / blue-dashed curve suggests a transmutation for $t_{\rm d}\leq t_{\rm Ga}$, whereas the left side of the red-solid / blue-dashed curve suggests that the transmutation may occur in the future. The dot-dashed brown and dashed magenta curves (equivalent to solid black curves of Fig. 5 and Fig. 7 of [60], respectively) are due to the constraints given by the LZ experiment for the nominal upper limit on the DM-nucleon spin-independent and DM-neutron spin-dependent cross section as a function of $m_{\chi}$, respectively. This figure shows the constrained $\sigma-m_{\chi}$ parameter spaces obtained due to the missing pulsar problem for the three different DM density profiles. See Sec. 7 for details.
@@
Now, let us consider a specific example with $\mathcal{D}\sim 10$ pc following [58, 59]. We will consider the Point (1) mentioned above, but not the Point (2), to be conservative. The results are shown in Fig. 4. The value of $\rho_{\rm DM}$ at a distance $\sim 10$ pc can be calculated to be $\sim 800$ GeV.cm-3 assuming the NFW DM density profile (Eq. 31). The parameter space located on the right side of the red-solid (blue-dashed) curve of Panel (a) of Fig. 4 indicates a transmutation for $t_{\rm d}<t_{\rm Ga}$. On the other hand, the left side of the red-solid (blue-dashed) curve of the same plot suggests that the transmutation may occur in the future. Thus, the Point (1) mentioned above and the DM density at $10$ pc give a lower limit of $\sigma$ as a function of $m_{\chi}$.
@@
The dot-dashed brown and dashed magenta curves in Panel (a) of Fig. 4 are due to the constraints given by the LUX-ZEPLIN (LZ 2022) experiment [60] for the nominal upper limit on the DM-nucleon spin-independent and DM-neutron spin-dependent cross section as a function of $m_{\chi}$, respectively. Note that the dot-dashed brown and dashed magenta curves in Panel (a) of Fig. 4 are equivalent to solid black curves of Fig. 5 and Fig. 7 of [60], respectively. Since both the LZ [60] and XENONnT [61] DM search experiments have considered the DM particles mass range upto $m_{\chi}\sim 10^{4}$ GeV, we indicate the upper limit of $\sigma$ upto the same $m_{\chi}$ value.
@@
Now, if the pulsars are missing from the region with distance $<10$ pc [58, 59] in the Galactic center, the $\sigma-m_{\chi}$ parameter space bounded by the red-solid curve and the dot-dashed brown curve of Panel (a) of Fig. 4 could be considered as the suitable constrained parameter space for the bosonic DM particles obtained from the missing pulsar problem assuming the DM-nucleon spin-independent interaction for the NFW DM density profile. Similarly, the $\sigma-m_{\chi}$ parameter space bounded by the red-solid curve and the dashed magenta curve of Panel (a) of Fig. 4 could be considered as the suitable constrained parameter space for bosonic DM particles obtained from the missing pulsar problem assuming the DM-neutron spin-dependent interaction. However, since the blue-dashed curve is located far from the dot-dashed brown and dashed magenta curves, the upper limit could not be suggested from the LZ and XENONnT experiments for the fermionic DM particles. Note that here we do not specifically consider a spin-independent or spin-dependent DM-nucleon interaction, but, by studying Panel (a), one can see that the possibility of the transmutation of neutron stars into BHs/naked singularities is more for the spin-dependent DM-neutron interaction.
@@ 8 Discussion and Conclusion
In this paper, we have shown that a dark core collapse could lead to not only a BH but also a naked singularity, depending on the initial condition of the DM cloud. A recent paper [53] has shown that some white dwarfs can realistically transmute into Kerr BHs and Kerr naked singularities ($a/M>1$) with a wide range of near- and sub-solar-mass values by capturing the asymmetric or non-self-annihilating DM particles. But that paper [53] has also shown that a neutron star, even if a rapidly spinning one, cannot transmute into a naked singularity with $a/M>1$ by accretion. On the other hand, the current paper, for the first time to the best of our knowledge, argues that a neutron star could transmute into a naked singularity.

### v2

@@ 7 Constraining $\sigma-m_{\chi}$ parameter space from the Missing pulsar problem
Point (1) should give a lower limit of $\sigma$ and Point (2) should give an upper limit of $\sigma$, each as a function of $m_{\chi},~{}\mathcal{D}$, DM profile, particle type (boson or fermion), etc. Thus, the $t_{\rm d}=t_{\rm Ga}$ condition should ideally give a value of $\sigma$ as a function of $m_{\chi},~{}\mathcal{D}$, DM profile, particle type (boson or fermion), etc. The transmutation timescale $t_{\rm d}$ can be calculated using eq. (5.3) and eq. (3.8).
(a)  NFW DM density profile (eq. 5.4): $\rho_{\rm DM}=800$ GeV.cm-3
(b)  Cored DM profile (eq. 7.1): $\rho_{\rm DM}=17$ GeV.cm-3
(c)  Cored DM profile (eq. 7.1): $\rho_{\rm DM}=410$ GeV.cm-3
@@
Figure 4: Constraints on the DM-nucleon scattering cross section $\sigma$, as a function of the PDM particle mass $m_{\chi}$ for the possible transmutation. The red-solid / blue-dashed curve corresponds to $t_{\rm d}=t_{\rm Ga}$ for boson/fermion, respectively, assuming all the old pulsars located within a distance $\mathcal{D}\sim 10$ pc around the Galactic center have been transmuted into BH or naked singularity. The parameter space located on the right side of the red-solid / blue-dashed curve suggests a transmutation for $t_{\rm d}\leq t_{\rm Ga}$, whereas the left side of the red-solid / blue-dashed curve suggests that the transmutation may occur in the future. The dot-dashed brown and dashed magenta curves (equivalent to solid black curves of figure 5 and figure 7 of [62], respectively) are due to the constraints given by the LZ experiment for the nominal upper limit on the DM-nucleon spin-independent and DM-neutron spin-dependent cross section as a function of $m_{\chi}$, respectively. This figure shows the constrained $\sigma-m_{\chi}$ parameter spaces obtained due to the missing pulsar problem for the three different DM density profiles. See section 7 for details.
@@
Now, let us consider a specific example with $\mathcal{D}\sim 10$ pc following [60, 61]. We will consider the Point (1) mentioned above, but not the Point (2), to be conservative. The results are shown in figure 4. The value of $\rho_{\rm DM}$ at a distance $\sim 10$ pc can be calculated to be $\sim 800$ GeV.cm-3 assuming the NFW DM density profile (eq. 5.4). The parameter space located on the right side of the red-solid (blue-dashed) curve of Panel (a) of figure 4 indicates a transmutation for $t_{\rm d}<t_{\rm Ga}$. On the other hand, the left side of the red-solid (blue-dashed) curve of the same plot suggests that the transmutation may occur in the future. Thus, the Point (1) mentioned above and the DM density at $10$ pc give a lower limit of $\sigma$ as a function of $m_{\chi}$.
@@
The dot-dashed brown and dashed magenta curves in Panel (a) of figure 4 are due to the constraints given by the LUX-ZEPLIN (LZ 2022) experiment [62] for the nominal upper limit on the DM-nucleon spin-independent and DM-neutron spin-dependent cross section as a function of $m_{\chi}$, respectively. Note that the dot-dashed brown and dashed magenta curves in Panel (a) of figure 4 are equivalent to solid black curves of figure 5 and figure 7 of [62], respectively. Since both the LZ [62] and XENONnT [63] DM search experiments have considered the DM particles mass range upto $m_{\chi}\sim 10^{4}$ GeV, we indicate the upper limit of $\sigma$ upto the same $m_{\chi}$ value.
@@
Now, if the pulsars are missing from the region with distance $<10$ pc [60, 61] in the Galactic center, the $\sigma-m_{\chi}$ parameter space bounded by the red-solid curve and the dot-dashed brown curve of Panel (a) of figure 4 could be considered as the suitable constrained parameter space for the bosonic DM particles obtained from the missing pulsar problem assuming the DM-nucleon spin-independent interaction for the NFW DM density profile. Similarly, the $\sigma-m_{\chi}$ parameter space bounded by the red-solid curve and the dashed magenta curve of Panel (a) of figure 4 could be considered as the suitable constrained parameter space for bosonic DM particles obtained from the missing pulsar problem assuming the DM-neutron spin-dependent interaction. However, since the blue-dashed curve is located far from the dot-dashed brown and dashed magenta curves, the upper limit could not be suggested from the LZ and XENONnT experiments for the fermionic DM particles. Note that here we do not specifically consider a spin-independent or spin-dependent DM-nucleon interaction, but, by studying Panel (a), one can see that the possibility of the transmutation of neutron stars into BHs/naked singularities is more for the spin-dependent DM-neutron interaction.
@@ 8 Discussion and Conclusion
In this paper, we have shown that a dark core collapse could lead to not only a BH but also a naked singularity, depending on the initial condition of the DM cloud. A recent paper [55] has shown that some white dwarfs can realistically transmute into Kerr BHs and Kerr naked singularities ($a/M>1$) with a wide range of near- and sub-solar-mass values by capturing the asymmetric or non-self-annihilating DM particles. But that paper [55] has also shown that a neutron star, even if a rapidly spinning one, cannot transmute into a naked singularity with $a/M>1$ by accretion. On the other hand, the current paper, for the first time to the best of our knowledge, argues that depending on the initial conditions, an endoparasitic naked singularity could form instead of an endoparasitic BH inside a neutron star, and the same neutron star could transmute into a naked singularity. Apart from different progenitors, there is a conceptual difference between the two papers. In [55], the naked singularity is proposed to form due to the accretion by an endoparasitic BH. This is a Kerr superspinar. Thus, a neutron star cannot give rise to such a naked singularity (see Appendix A of [55]), but a white dwarf can. Consequently, the mass of the naked singularity cannot be more than that of a white dwarf (implying mostly sub-solar or slightly super-solar mass). On the other hand, in the current paper, a naked singularity is formed directly by collapse and not from a BH. It is not a Kerr superspinar and hence a different kind of naked singularity from the former. This is the main conceptual difference. In this case, compared to the previous one, usually mass should be more but the spin should be less, when the collapse happens at the core of a neutron star. Even if the end product is a BH, and not a naked singularity, the mass of the final BH for the former paper (i.e., [55]) should typically be less than that for the current paper. These differences could also be helpful to determine whether the progenitor was a neutron star or a white dwarf. However, it is possible that a low-mass BH or a low-mass naked singularity with a low spin parameter value could form via the accretion by an endoparasitic collapsed object within either a white dwarf or a neutron star of similar masses and similarly low spin parameter values. In such a case, the progenitor cannot be easily identified either as a white dwarf or as a neutron star. While a detailed study of the said accretion process might provide a way to identify the progenitor in this case, such a process is not yet well-studied, would involve detailed fluid dynamical computations, and is not within the scope of this paper.
@@
In this paper, we have considered the non-self-annihilating DM particles with a mass range: $1$ GeV $<m_{\chi}<10^{6}$ GeV [28, 55]. This is sufficient for our purpose of studying the formation of low mass collapsed objects which could have been involved in GW190425 and GW190814. Our result is also commensurate with both the LZ [62] and XENONnT [63] DM search experiments, as seen from figure 4. It is hardly possible to form such a low mass collapsed object by accumulating the DM particles of mass less than the MeV range within the time $t_{\rm Ga}$, as seen from figure 2. However, in a recent paper [69], the supermassive black hole formation has been proposed from the gravitational collapse of fermionic dense DM cores of particles of keV mass range at the centre of DM haloes.